OpenStudy (ellerai):

Dave's parents give him at least $70 but not more than $120 each month for public transportation. If he takes the train, he has to pay a fee once a month of $10 and then $6.50 each day for a ticket. What is the greatest number of days that Dave can take the train each month and still remain within his public transportation budget?

OpenStudy (cas_fangirl_14):

This is going to take a minute to write out.

OpenStudy (ellerai):

Okay! :)

OpenStudy (cas_fangirl_14):

If Dave gets $120. $120-$10=$110 then 110/6.50=16.923 Dave can take the train 16 times.

OpenStudy (ellerai):

Thank you! can you help me with one more?

OpenStudy (cas_fangirl_14):

Sure :)

OpenStudy (ellerai):

Choose the equation you would use to solve the problem. It takes Zane m minutes to clean his room. This is 35 minutes more than the 22 minutes it takes him to walk the dog. How long does it take him to clean his room? A. m + 22 = 35 B. m - 35 = 22 C. 35 - m = 22 D. m + 35 = 22

OpenStudy (cas_fangirl_14):

It's B The problem tells you it takes 35 and 22 min to clean his room. You just have to rearrange the terms. :)

OpenStudy (ellerai):

thank you so much!!

OpenStudy (cas_fangirl_14):

\(\color{Purple}{No~Problem~~~:)}\)